# Convergence of empirical Gromov-Wasserstein distance

### v1

@@ Abstract.
We study rates of convergence for estimation of the Gromov-Wasserstein distance. For two marginals supported on compact subsets of $\mathbb{R}^{d_{x}}$ and $\mathbb{R}^{d_{y}}$, respectively, with $\min\{d_{x},d_{y}\}>4$, prior work established the rate $n^{-\frac{2}{\min\{d_{x},d_{y}\}}}$ for the plug-in empirical estimator based on $n$ i.i.d. samples. We extend this fundamental result to marginals with unbounded supports, assuming only finite polynomial moments. Our proof techniques for the upper bounds can be adapted to obtain sample complexity results for penalized Wasserstein alignment that encompasses the Gromov-Wasserstein distance and Wasserstein Procrustes in unbounded settings. Furthermore, we establish matching minimax lower bounds (up to logarithmic factors) for estimating the Gromov-Wasserstein distance.
@@
K. Kato is partially supported by NSF grant DMS-2413405.
@@ 1.1. Overview
where $\Pi(\mu,\nu)$ denotes the set of couplings for $\mu$ and $\nu$.
@@
The recent work [ZGMS24] established the first sample complexity results for the GW distance with $(p,q)=(2,2)$ by leveraging a variational representation of $\mathsf{GW}_{2,2}^{2}$ that links the GW problem to standard OT. Using this representation and OT duality theory, they showed that for compactly supported $\mu$ and $\nu$, the empirical estimator based on $n$ i.i.d. samples converges at the rate $n^{-2/(d_{x}\wedge d_{y})}$ with $d_{x}\wedge d_{y}=\min\{d_{x},d_{y}\}$ when $d_{X}\wedge d_{y}>4$. [^1] The empirical estimator is defined by plugging in the empirical distributions for $\mu$ and $\nu$. Notably, this rate adapts automatically to the smaller of the two dimensions, rather than being governed by the worst case dimension $d_{x}\vee d_{y}=\max\{d_{x},d_{y}\}$. This is achieved by adapting the lower complexity adaptation (LCA) principle studied in [HSM24] in the OT case. They further derived matching lower bounds for the empirical estimator, but did not derive minimax lower bounds. As such, strictly speaking, their lower bound result does not rule out the possibility that estimators other than the plug-in type could uniformly outperform the empirical estimator.
The first goal of this work is to extend their fundamental sample complexity result to the unbounded support case. Our result establishes that, when $d_{x}\wedge d_{y}>4$, the $n^{-2/(d_{x}\wedge d_{y})}$ rate (without logarithmic factors) continues to hold for the empirical GW even when $\mu,\nu$ only possess finite polynomial moments. It is worth noting that, even in OT cost estimation, extending results from the compact support case to the unbounded setting is often highly nontrivial. This is because global regularity estimates for dual potentials, which are often available for the compact support case, do not continue to hold for the unbounded setting, and establishing local regularity estimates would require delicate tail conditions on the marginals; see [CF21, MBNWW24] and the discussion in Section 2. Our proof essentially builds on an adaptation of an idea in [SH25], but with some nontrivial twists; see the discussion after Theorem 3.1 below for details. Furthermore, our proof technique can be adapted to obtain sample complexity results for penalized Wasserstein alignment [PSW25] that encompasses the Gromov-Wasserstein distance and Wasserstein Procrustes in unbounded settings.
@@ 1.2. Literature review
Convergence and exact asymptotics of empirical OT costs have been extensively studied in the statistics and probability literature; see, e.g., [Dud69, AKT84, Tal92, Tal94, DY95, BdMM02, BB13, DSS13, BLG14, FG15, WB19, Lei20, CRL+20, MNW24, SH25]. Most of these references focus on establishing sharp rates for empirical distributions under $p$-Wasserstein distances $W_{p}$ with $p\in[1,\infty)$. [^2] See Section 2 for the definition of $W_{p}$. For instance, let $\mu$ be a Borel probability measure on $\mathbb{R}^{d}$ and $\hat{\mu}_{n}$ be the empirical distribution for $n$ i.i.d. samples from $\mu$; then, [FG15] showed that when $\mu$ has a finite $q$-th moment with $q>p$,
@@
where $\lesssim$ denotes an inequality holding up to a numerical constant that is independent of $n$ but may depend on other parameters. As a canonical case, when $d>4$ and q>2d/(d2)q>2d/(d-2), the result (2) implies $\mathbb{E}\big{[}W_{2}^{2}(\hat{\mu}_{n},\mu)\big{]}\lesssim n^{-2/d}$. These rates are known to be sharp in various settings with a notable exception of the $d=2p$ case; cf. the discussion after Theorem 1 in [FG15].
@@ 1.3. Organization
The rest of the paper is organized as follows. Section 2 collects brief overviews of OT and GW and a discussion on the prior sample complexity results for GW. Section 3 presents the main results on sample complexity in the unbounded setting and minimax lower bounds for GW. All proofs are gathered in Section 4.
@@ 2.2. Gromov-Wasserstein distance
    $\mathsf{GW}_{p,q}(\mu,\nu)\leq q\,2^{q+1-\frac{1}{p}+\frac{1}{pq}}\big{(}m_{pq}(\mu)+m_{pq}(\nu)\big{)}^{\frac{q-1}{pq}}W_{pq}(\mu,\nu).$
@@
where $c_{A}(x,y):=-4\|x\|^{2}\|y\|^{2}-32x^{\top}Ay$. Furthermore, the infimum on the right-hand side is achieved by some $A\in\mathbb{R}^{d_{x}\times d_{y}}$ with $\|A\|_{\operatorname{op}}\leq\sqrt{m_{2}(\mu)m_{2}(\nu)}/2$.
@@
with equality holding if and only if $A=\frac{1}{2}\int xy^{\top}\,d\pi(x,y)$, where $\langle\cdot,\cdot\rangle_{\mathrm{F}}$ denotes the Frobenius inner product and $\langle A,\int xy^{\top}\,d\pi(x,y)\rangle_{\mathrm{F}}=\int x^{\top}Ay\,d\pi(x,y)$. Interchanging $\inf_{\pi}$ and $\inf_{A}$ gives the expression (6). Finally, since $\|\int xy^{\top}\,d\pi(x,y)\|_{\operatorname{op}}\leq\sqrt{m_{2}(\mu)m_{2}(\nu)}$, $\inf_{A}$ in (6) can be reduced to the infimum over $A$ with $\|A\|_{\operatorname{op}}\leq\sqrt{m_{2}(\mu)m_{2}(\nu)}/2$. As the mapping $A\mapsto 32\|A\|_{\mathrm{F}}^{2}+T_{c_{A}}(\mu,\nu)$ is continuous, the final claim follows.
@@ 2.3. Prior results for upper bounds and challenges in unbounded settings
$\varphi_{n,m}:=(n\wedge m)^{-\frac{2}{(d_{x}\wedge d_{y})\vee 4}}(\log(n\vee m))^{\mathbbm{1}_{\{d_{x}\wedge d_{y}=4\}}}.$ (8)
@@ 3.1. Upper bounds for marginals with unbounded supports
$\sup_{\begin{subarray}{c}(\mu,\nu)\in\mathcal{P}_{4q}(\mathcal{X})\times\mathcal{P}_{4q}(\mathcal{Y})\\ m_{4q}(\mu)\vee m_{4q}(\nu)\leq M\end{subarray}}\mathbb{E}\left[\big{|}D(\hat{\mu}_{n},\hat{\nu}_{m})-D(\mu,\nu)\big{|}\right]\lesssim\varphi_{n,m}+(n\wedge m)^{-\frac{1}{2}}\sqrt{\log(n\wedge m)}\\ +(n\wedge m)^{\frac{1-q}{q}}\Big{\{}(n\wedge m)^{\frac{2(d_{x}\vee d_{y})}{q}}\wedge(n\wedge m)^{\frac{d_{x}d_{y}}{2q}}\Big{\}}\log(n\wedge m),$ (11)
where the hidden constant depends only on $q,d_{x},d_{y}$ and $M$.
@@
the empirical estimator $D(\hat{\mu}_{n},\hat{\nu}_{m})$ achieves the rate $(n\wedge m)^{-\frac{2}{d_{x}\wedge d_{y}}}(\log(n\vee m))^{\mathbbm{1}_{\{d_{x}\wedge d_{y}=4\}}}$ that was known to hold only for compactly supported marginals. The minimal moment condition (12) scales linearly in $d_{x}\vee d_{y}$.
@@
$\sup_{\begin{subarray}{c}(\mu,\nu)\in\mathcal{P}_{8\vee(2q)}(\mathcal{X})\times\mathcal{P}(B_{\mathcal{Y}}(0,1))\\ m_{8\vee(2q)}(\mu)\leq M\end{subarray}}\mathbb{E}\left[\big{|}D(\hat{\mu}_{n},\hat{\nu}_{m})-D(\mu,\nu)\big{|}\right]\lesssim\bar{\varphi}_{n,m,q},$
where the hidden constant depends only on $q,d_{x},d_{y}$ and $M$.
The proposition requires $\mu$ to have a finite $\big{(}8\vee(2q)\big{)}$-th moment instead of $4q$, while implying that the empirical estimator retains the rate $n^{-\frac{2}{d_{x}\vee d_{y}}}(\log(n\vee m))^{\mathbbm{1}_{\{d_{x}\wedge d_{y}=4\}}}$ when $d_{x}\wedge d_{y}\geq 4$ provided that $q$ satisfies (12). The reason behind this is that, in Theorem 3.1 the cost $c_{A}$ is bounded as $|c_{A}(x,y)|\lesssim 1+\|x\|^{4}+\|y\|^{4}$, while if $\nu$ is supported in $B_{\mathcal{Y}}(0,1)$, one can use a bound $|c_{A}(x,y)|\lesssim 1+\|x\|^{2}$. This implies that dual potentials for $c_{A}$ have bounded $q$-th moments if $\mu$ has a finite $(2q)$-th moment. The proof of the proposition is a reasonably minor modification of that for Theorem 3.1. In principle, one can also consider more general situations where two marginals have finite moments of different orders, which, however, will not be pursued here.
@@
where $\mathcal{O}(d)$ denotes the set of $d\times d$ orthogonal matrices and $\tilde{c}_{O}(x,y)=\|Ox-y\|^{2}=\|x\|^{2}+\|y\|^{2}-2x^{\top}Oy$.
@@
$\mathbb{E}\big{[}|\tilde{W}_{2}^{2}(\hat{\mu}_{n},\hat{\nu}_{m})-\tilde{W}_{2}^{2}(\mu,\nu)|\big{]}\lesssim(n\wedge m)^{-\frac{2}{d}}(\log(n\vee m))^{\mathbbm{1}_{\{d=4\}}}+(n\wedge m)^{-\frac{1}{2}}\sqrt{\log(n\wedge m)}\\ +(n\wedge m)^{\frac{1-q}{q}}\Big{\{}(n\wedge m)^{\frac{2d}{q}}\wedge(n\wedge m)^{\frac{d^{2}}{2q}}\Big{\}}\log(n\wedge m).$
The hidden constant depends only on $q,d,m_{2q}(\mu)$ and $m_{2q}(\nu)$.
The proof is almost identical to Theorem 3.1, given that a version of Lemma 4.2 below (with $r^{4}$ replaced by $r^{2}$) holds for $T_{\tilde{c}_{O}}$ (see [SH25]) and hence omitted for brevity. Furthermore, a minor modification of the proof of Theorem 3.2 below yields that $(n\wedge m)^{-\frac{2}{d}}$ matches a minimax lower bound (up to a logarithmic factor) over the class of distributions supported in the unit ball.

### v2

@@ Abstract.
We study rates of convergence for estimation of the Gromov-Wasserstein (GW) distance. For two marginals supported on compact subsets of $\mathbb{R}^{d_{x}}$ and $\mathbb{R}^{d_{y}}$, respectively, with $\min\{d_{x},d_{y}\}>4$, prior work established the rate $n^{-\frac{2}{\min\{d_{x},d_{y}\}}}$ in $L^{1}$ for the plug-in empirical estimator based on $n$ i.i.d. samples. We extend this fundamental result to marginals with unbounded supports, assuming only finite polynomial moments. Our proof techniques for the upper bounds can be adapted to obtain sample complexity results for penalized Wasserstein alignment that encompasses the GW distance and Wasserstein Procrustes. Furthermore, we establish matching minimax lower bounds (up to logarithmic factors) for estimating the GW distance. Finally, we establish deviation inequalities for the error of empirical GW in cases where two marginals have compact supports, exponential tails, or finite polynomial moments. The deviation inequalities yield that the same rate $n^{-\frac{2}{\min\{d_{x},d_{y}\}}}$ holds for empirical GW also with high probability.
@@
K. Kato is partially supported by NSF grant DMS-2413405. We would like to thank Ziv Goldfeld and Sloan Nietert for stimulating discussions.
@@ 1.1. Overview
where $\Pi(\mu,\nu)$ denotes the set of couplings for $\mu$ and $\nu$. Recall that any coupling $\pi\in\Pi(\mu,\nu)$ is a joint distribution on $\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}$ with marginals $\mu,\nu$ on $\mathbb{R}^{d_{x}},\mathbb{R}^{d_{y}}$, respectively.
@@
The recent work [ZGMS24] established the first sample complexity results for the GW distance with $(p,q)=(2,2)$ by leveraging a variational representation of $\mathsf{GW}_{2,2}^{2}$ that links the GW problem to standard OT. Using this representation and OT duality theory, they showed that for compactly supported $\mu$ and $\nu$, the empirical estimator based on $n$ i.i.d. samples converges, in $L^{1}$, at the rate $n^{-2/(d_{x}\wedge d_{y})}$ with $d_{x}\wedge d_{y}=\min\{d_{x},d_{y}\}$ when $d_{X}\wedge d_{y}>4$. [^1] The empirical estimator is defined by plugging in the empirical distributions for $\mu$ and $\nu$. Notably, this rate adapts automatically to the smaller of the two dimensions, rather than being governed by the worst case dimension $d_{x}\vee d_{y}=\max\{d_{x},d_{y}\}$. This is achieved by adapting the lower complexity adaptation (LCA) principle studied in [HSM24] in the OT case. They further derived matching lower bounds for the empirical estimator, but did not derive minimax lower bounds. As such, strictly speaking, their lower bound result does not rule out the possibility that estimators other than the plug-in type could uniformly outperform the empirical estimator.
The first goal of this work is to extend their fundamental sample complexity result to the unbounded support case. Our result establishes that, when $d_{x}\wedge d_{y}>4$, the $n^{-2/(d_{x}\wedge d_{y})}$ rate (without logarithmic factors) continues to hold for the empirical GW even when $\mu,\nu$ only possess finite polynomial moments. It is worth noting that, even in OT cost estimation, extending results from the compact support case to the unbounded setting is often highly nontrivial. This is because global regularity estimates for dual potentials, which are often available for the compact support case, do not continue to hold for the unbounded setting, and establishing local regularity estimates would require delicate tail conditions on the marginals; see [CF21, MBNWW24] and the discussion in Section 2. Our proof essentially builds on an adaptation of an idea in [SH25], but with some nontrivial twists; see the discussion after Theorem 3.1 below for details. Furthermore, our proof technique can be adapted to obtain sample complexity results for penalized Wasserstein alignment [PSW25] that encompasses the GW distance and Wasserstein Procrustes.
In addition, we establish two auxiliary results. First, when one of the marginals is heavy-tailed with less than 8-th moments, we show that the rate of convergence of the empirical GW distance can be arbitrarily slow. The result sheds new light on the tradeoff between heavy-tailedness of the distributions and the speed of convergence of the empirical distance. Second, in the semidiscrete setting, i.e., when one of the marginals, say $\nu$, is finitely discrete, we show that the parametric convergence rate $n^{-1/2}$ holds whenever the other marginal $\mu$ has a finite 8-th moment. The result complements a recent limiting distributional result for semidiscrete GW in [RGK24b] with compactly supported $\mu$ (cf. Theorem 7 there).
@@
Our third goal is to establish deviation inequalities for the error of empirical GW, or more precisely, the discrepancy between the squared empirical and population GW distances. We consider the three scenarios where two marginals have (i) compact supports, (ii) exponential tails (more precisely, finite $\psi_{\beta}$-norms for some $\beta>0$), and (iii) finite polynomial moments. For the first two cases, exponential deviation holds, while for the last case, only polynomial deviation holds. Our result shows that, when $d_{x}\wedge d_{y}>4$, the discrepancy between the squared empirical and population GW distances is at most a constant multiple of $n^{-2/(d_{x}\wedge d_{y})}$ with high probability, complementing the upper bounds in expectation. Our proof for the deviation inequalities for the unbounded support cases builds on a variant of McDiarmid’s inequality due to [Com24] that only requires the bounded difference condition to hold on a high-probability event. The deviation inequalities for empirical GW are new, even for the compactly supported case.
In sum, this work establishes novel sample complexity upper bounds and deviation inequalities for empirical GW in possibly unbounded settings and derives minimax lower bounds for GW estimation. These results close several important gaps in the literature and contribute to a deeper understanding of the GW estimation problem.
@@ 1.2. Literature review
Convergence and exact asymptotics of empirical OT costs have been extensively studied in the statistics and probability literature; see, e.g., [AKT84, Tal92, Tal94, DY95, dBGM99, BdMM02, BB13, DSS13, BLG14, FG15, WB19, Lei20, CRL+20, MNW24, SH25]. Most of these references focus on establishing sharp rates for empirical distributions under $p$-Wasserstein distances $W_{p}$ with $p\in[1,\infty)$. [^2] See Section 2 for the definition of $W_{p}$. For instance, let $\mu$ be a Borel probability measure on $\mathbb{R}^{d}$ and $\hat{\mu}_{n}$ be the empirical distribution for $n$ i.i.d. samples from $\mu$; then, [FG15] showed that when $\mu$ has a finite $q$-th moment with $q>p$,
@@
where $\lesssim$ denotes an inequality holding up to a numerical constant that is independent of $n$ but may depend on other parameters. As a canonical case, when $d>4$ and q>2d/(d2)q>2d/(d-2), (2) implies $\mathbb{E}\big{[}W_{2}^{2}(\hat{\mu}_{n},\mu)\big{]}\lesssim n^{-2/d}$. These rates in (2) are known to be sharp in various settings with a notable exception of the $d=2p$ case; cf. the discussion after Theorem 1 in [FG15]. The study of the empirical OT cost for heavy-tailed marginals is relatively scarce, to the best of the authors’ knowledge. One exception is [dBGM99], where the authors established limit theorems and moment convergence of $W_{1}(\hat{\mu}_{n},\mu)$ in $d=1$ when $\mu$ is in the domain of attraction of an $\alpha$-stable law with $\alpha\in(1,2]$.
@@
Concentration or deviation inequalities for empirical OT costs, akin to our Theorem 3.3, seem to have been less explored. One related result is Theorem 2 in [CRL+20] that establishes $\mathbb{P}(|W_{2}^{2}(\hat{\mu}_{n},\hat{\nu}_{n})-\mathbb{E}[W_{2}^{2}(\hat{\mu}_{n},\hat{\nu}_{n})]|\geq t)\leq 2e^{-nt^{2}}$, when $\mu,\nu$ are supported in a set of diameter $1$. Combining (3), the preceding result yields a deviation inequality for $|W_{2}^{2}(\hat{\mu}_{n},\hat{\nu}_{n})-W_{2}^{2}(\mu,\nu)|$. Beyond the compact support case, both [MNW24, SH25] did not study concentration or deviation inequalities for errors of empirical OT costs when $\mu\neq\nu$. Of note is that one can use the decomposition $|W_{p}(\hat{\mu}_{n},\hat{\nu}_{m})-W_{p}(\mu,\nu)|\leq W_{p}(\hat{\mu}_{n},\mu)+W_{p}(\hat{\nu}_{m},\nu)$ and apply known concentration or deviation inequalities for $W_{p}(\hat{\mu}_{n},\mu)$ and $W_{p}(\hat{\nu}_{m},\nu)$, e.g., in [FG15], to obtain deviation inequalities for $|W_{p}(\hat{\mu}_{n},\hat{\nu}_{m})-W_{p}(\mu,\nu)|$, but the resulting inequalities are suboptimal because $|W_{p}(\hat{\mu}_{n},\hat{\nu}_{m})-W_{p}(\mu,\nu)|$ should scale faster than $\max\{W_{p}(\hat{\mu}_{n},\mu),W_{p}(\hat{\nu}_{m},\nu)\}$ when $\mu\neq\nu$; cf. the discussion below (3).
@@ 1.3. Organization
The rest of the paper is organized as follows. Section 2 collects brief overviews of OT and GW and a discussion on the prior sample complexity results for GW. Section 3 presents the main results on sample complexity in the unbounded setting, minimax lower bounds, and deviation inequalities for GW. All proofs are gathered in Section 4.
@@ 2.2. Gromov-Wasserstein distance
    $\mathsf{GW}_{p,q}(\mu,\nu)\leq q\,2^{q+1-\frac{1}{p}+\frac{1}{pq}}\big{(}\mathfrak{m}_{pq}(\mu)+\mathfrak{m}_{pq}(\nu)\big{)}^{\frac{q-1}{pq}}W_{pq}(\mu,\nu).$
@@
where $c_{A}(x,y):=-4\|x\|^{2}\|y\|^{2}-32x^{\top}Ay$. Furthermore, the infimum on the right-hand side is achieved by some $A\in\mathbb{R}^{d_{x}\times d_{y}}$ with $\|A\|_{\operatorname{op}}\leq\sqrt{\mathfrak{m}_{2}(\mu)\mathfrak{m}_{2}(\nu)}/2$.
@@
with equality holding if and only if $A=\frac{1}{2}\int xy^{\top}\,d\pi(x,y)$, where $\langle\cdot,\cdot\rangle_{\mathrm{F}}$ denotes the Frobenius inner product and $\langle A,\int xy^{\top}\,d\pi(x,y)\rangle_{\mathrm{F}}=\int x^{\top}Ay\,d\pi(x,y)$. Interchanging $\inf_{\pi}$ and $\inf_{A}$ gives the expression (6). Finally, since $\|\int xy^{\top}\,d\pi(x,y)\|_{\operatorname{op}}\leq\sqrt{\mathfrak{m}_{2}(\mu)\mathfrak{m}_{2}(\nu)}$, $\inf_{A}$ in (6) can be reduced to the infimum over $A$ with $\|A\|_{\operatorname{op}}\leq\sqrt{\mathfrak{m}_{2}(\mu)\mathfrak{m}_{2}(\nu)}/2$. As the mapping $A\mapsto 32\|A\|_{\mathrm{F}}^{2}+T_{c_{A}}(\mu,\nu)$ is continuous, the final claim follows.
@@ 2.3. Prior results for upper bounds and challenges in unbounded settings
$\varphi_{n,m}:=(n\wedge m)^{-\frac{2}{(d_{x}\wedge d_{y})\vee 4}}\big{(}\log(n\wedge m)\big{)}^{\mathbbm{1}_{\{d_{x}\wedge d_{y}=4\}}}.$ (8)
@@ 3.1. Upper bounds for marginals with unbounded supports
$\sup_{\begin{subarray}{c}(\mu,\nu)\in\mathcal{P}_{4q}(\mathcal{X})\times\mathcal{P}_{4q}(\mathcal{Y})\\ \mathfrak{m}_{4q}(\mu)\vee\mathfrak{m}_{4q}(\nu)\leq M\end{subarray}}\mathbb{E}\left[\big{|}D(\hat{\mu}_{n},\hat{\nu}_{m})-D(\mu,\nu)\big{|}\right]\lesssim\varphi_{n,m}+(n\wedge m)^{-\frac{1}{2}}\sqrt{\log(n\wedge m)}\\ +(n\wedge m)^{\frac{1-q}{q}}\Big{\{}(n\wedge m)^{\frac{2(d_{x}\vee d_{y})}{q}}\wedge(n\wedge m)^{\frac{d_{x}d_{y}}{2q}}\Big{\}}\log(n\wedge m),$ (11)
where the hidden constant depends only on $d_{x},d_{y},q$ and $M$.
@@
the empirical estimator $D(\hat{\mu}_{n},\hat{\nu}_{m})$ achieves the rate $\varphi_{n,m}$ that was known to hold only for compactly supported marginals. The minimal moment condition (12) scales linearly in $d_{x}\vee d_{y}$.
@@
$\sup_{\begin{subarray}{c}(\mu,\nu)\in\mathcal{P}_{2q}(\mathcal{X})\times\mathcal{P}(B_{\mathcal{Y}}(0,1))\\ \mathfrak{m}_{2q}(\mu)\leq M\end{subarray}}\mathbb{E}\left[\big{|}D(\hat{\mu}_{n},\hat{\nu}_{m})-D(\mu,\nu)\big{|}\right]\lesssim\bar{\varphi}_{n,m,q}+n^{\frac{2-q}{q}},$
where the hidden constant depends only on $d_{x},d_{y},q$ and $M$.
The proposition requires $\mu$ to have a finite $(2q)$-th moment instead of $4q$, while implying that the empirical estimator retains the rate $\varphi_{n,m}$ when $d_{x}\wedge d_{y}\geq 4$ provided that $q$ satisfies (12) (which entails $q>4$ so that $n^{\frac{2-q}{q}}=o(n^{-1/2})$). The reason behind this is that, in Theorem 3.1 the cost $c_{A}$ is bounded as $|c_{A}(x,y)|\lesssim 1+\|x\|^{4}+\|y\|^{4}$, while if $\nu$ is supported in $B_{\mathcal{Y}}(0,1)$, one can use a bound $|c_{A}(x,y)|\lesssim 1+\|x\|^{2}$. This implies that dual potentials for $c_{A}$ have bounded $q$-th moments if $\mu$ has a finite $(2q)$-th moment. The proof of the proposition is a reasonably minor modification of that for Theorem 3.1. In principle, one can also consider more general situations where two marginals have finite moments of different orders, which, however, will not be pursued here.
The extra $n^{\frac{2-q}{q}}$ factor comes from the fact that $S_{1}(\hat{\mu}_{n},\hat{\nu}_{m})-S_{1}(\mu,\nu)$ can be approximated as $\frac{2}{n}\sum_{i=1}^{n}(\|X_{i}\|^{4}-\mathfrak{m}_{4}(\mu))$ and that $\|X_{i}\|^{4}$ has only a finite $(q/2)$-th moment. In fact, if $\mu$ is heavy-tailed and has less than 8-th moments, the rate of convergence for $D(\hat{\mu}_{n},\hat{\nu}_{m})$ can be arbitrarily slow, as the next proposition demonstrates.
Proposition 3.2 (Heavy-tailed case).
Let $d_{x}=d_{y}=1,n=m$, and $\nu\in\mathcal{P}(B_{\mathbb{R}}(0,1))$. For any $\alpha\in(1,2)$, there exists a distribution $\mu$ on $\mathbb{R}$ for which the following holds: for $X\sim\mu$,
$\mathbb{E}[|X|^{4\alpha}]=\infty,\quad\mathbb{E}[|X|^{2q}]<\infty\ \text{for}\ q\in(0,2\alpha),$ (13)
and
$\liminf_{n\to\infty}n^{\frac{\alpha-1}{\alpha}}\mathbb{E}\left[\big{|}D(\hat{\mu}_{n},\hat{\nu}_{n})-D(\mu,\nu)\big{|}\right]>0.$ (14)
The proof constructs $\mu$ so that $X^{4}$ with $X\sim\mu$ is in the domain of attraction of an $\alpha$-stable law. [^4] See Chapter 9 in [Bre92] for stable distributions. Indeed, the proof derives an explicit lower bound
$\mathbb{E}\left[\big{|}D(\hat{\mu}_{n},\hat{\nu}_{n})-D(\mu,\nu)\big{|}\right]\geq n^{\frac{1-\alpha}{\alpha}}\mathbb{E}[|Z_{\alpha}|]+o\big{(}n^{\frac{1-\alpha}{\alpha}}\big{)},$
where $Z_{\alpha}$ follows a symmetric $\alpha$-stable law with characteristic function $e^{-\mathsf{c}_{\alpha}|t|^{\alpha}}$ with $\mathsf{c}_{\alpha}:=\alpha\int_{0}^{\infty}(1-\cos x)x^{-\alpha-1}\,dx$.
Finally, as another special case, we shall consider the semidiscrete case where $\mu$ is general but $\nu$ is finitely discrete, so that $\nu$ is of the form
$\nu=\sum_{j=1}^{\ell}\nu_{j}\delta_{y_{j}},$
where $(\nu_{1},\dots,\nu_{\ell})^{\top}$ is a simplex vector with positive elements (i.e., $\nu_{j}>0$ for all $j$ and $\sum_{j=1}^{\ell}\nu_{j}=1$) and $\{y_{1},\dots,y_{\ell}\}\subset\mathcal{Y}$. For simplicity, we assume $\max_{1\leq j\leq\ell}\|y_{j}\|\leq 1$.
Proposition 3.3 (Semidiscrete case).
Consider the semidiscrete setting considered above. For given $q\in(2,\infty)$ and $M\geq 1$, we have
$\sup_{\mu\in\mathcal{P}_{2q}(\mathcal{X}):\mathfrak{m}_{2q}(\mu)\leq M}\mathbb{E}\left[\big{|}D(\hat{\mu}_{n},\hat{\nu}_{m})-D(\mu,\nu)\big{|}\right]\lesssim(n\wedge m)^{-1/2}+n^{\frac{2-q}{q}},$
where the hidden constant depends only on $d_{x},d_{y},q,M,\ell$ and $\min_{1\leq j\leq\ell}\nu_{j}$.
In particular, if $\mu$ has a finite 8-th moment,
$\mathbb{E}\left[\big{|}D(\hat{\mu}_{n},\hat{\nu}_{m})-D(\mu,\nu)\big{|}\right]\lesssim(n\wedge m)^{-1/2}.$
In contrast to Theorem 3.1 and Proposition 3.1, the bound in Proposition 3.3 does not involve the extra $\sqrt{\log(n\wedge m)}$ factor. This is because, in the semidiscrete setting, the dual potentials $(f_{A},g_{A})$ have simple expressions that enable us to avoid using discretization of the set of $A$ (cf. the discussion after Theorem 3.1).
@@
where $\mathcal{O}(d)$ denotes the set of $d\times d$ orthogonal matrices and $\tilde{c}_{O}(x,y)=\|Ox-y\|^{2}=\|x\|^{2}+\|y\|^{2}-2x^{\top}O^{\top}y$.
@@
$\mathbb{E}\big{[}\big{|}\tilde{W}_{2}^{2}(\hat{\mu}_{n},\hat{\nu}_{m})-\tilde{W}_{2}^{2}(\mu,\nu)\big{|}\big{]}\lesssim(n\wedge m)^{-\frac{2}{d}}(\log(n\wedge m))^{\mathbbm{1}_{\{d=4\}}}+(n\wedge m)^{-\frac{1}{2}}\sqrt{\log(n\wedge m)}\\ +(n\wedge m)^{\frac{1-q}{q}}\Big{\{}(n\wedge m)^{\frac{2(d-1)}{q}}\wedge(n\wedge m)^{\frac{d(d-1)}{2q}}\Big{\}}\log(n\wedge m).$
The hidden constant depends only on $q,d,\mathfrak{m}_{2q}(\mu)$ and $\mathfrak{m}_{2q}(\nu)$.
The proof is almost identical to Theorem 3.1, given that a version of Lemma 4.2 below (with $r^{4}$ replaced by $r^{2}$) holds for $T_{\tilde{c}_{O}}$ (see [SH25]) and that the $\epsilon$-covering number of the orthogonal group $\mathcal{O}(d)$ under $\|\cdot\|_{\operatorname{op}}$ is at most $(K/\epsilon)^{d(d-1)/2}$ for $0<\epsilon<1$ for some universal constant $K$ (cf. Theorem 7 in [Sza98]). Since the modification is minor, we omit the proof for brevity. Furthermore, a minor modification of the proof of Theorem 3.2 below yields that $(n\wedge m)^{-\frac{2}{d\vee 4}}$ matches a minimax lower bound (up to a logarithmic factor) over the class of distributions supported in the unit ball.